Find the quotient when the difference of 985 and 958 is divided by 9.

Grade:8

2 Answers

Saurabh Koranglekar
3 years ago
The difference between the numbers 985 and 958 when divided by 9, we know that when ten’s and unit’s place is interchanged we get quotient as a difference of unit’s and ten’s place.

∴Quotient is 8 – 5 = 3
Vikas TU
3 years ago
The difference between the numbers 985 and 958 when divided by 9, we know that when ten's and unit's place is interchanged we get quotient as a difference of unit's and ten's place.
985 - 958 = 27
27÷ 9 = 3